OpenStudy (anonymous):

Consider the following pair of equations: y = −2x + 8 y = x − 1 Explain how you will solve the pair of equations by substitution. Show all the steps and write the solution in (x, y) form. please help explain idk :(

OpenStudy (anonymous):

Substitute y= x-1 into y -->> y= -2x + 8 x-1= -2x + 8 Add 1 both sides and +2x which then gives you 3x= 9 the divide by 3 and your answer should be x= 3

OpenStudy (anonymous):

how would I find y? I already got x but the y part confuses me ;/

OpenStudy (er.mohd.amir):

both y is same so -2x+8=x-1 -2x-x=-8-1 -3x=-9 3x=9 x=9/3 x=3 y=x-1 y=3-1 y=2

OpenStudy (anonymous):

Thank you I really appreciate it. This helped me understand how to do this :).